# A comprehensive study of the relations between the properties of planetary systems and the chemical compositions of their host stars

### v1

@@ IV.4.2 Solar twins
Focusing now on host stars with parameters very similar to the solar values, previous works in the literature have found that the Sun is depleted in refractory elements relative to solar twins (e.g., melendez09; bedell18; rampalli24; sun25a; sun25b). If we assume here that solar twins are stars that have stellar parameters within the ranges 5677 K 5877 K, 4.34 4.54 and -0.10 dex [Fe/H] 0.10 dex, there are 25 solar twins in our sample. Among these systems, we have 6 SE, 11 SN, 6 SE + SN, 1 SN + SS and 1 JP.
@@
It is worth noting that four of these stars (KOI-444, KOI-3232, KOI-4383 and KOI-4400) have ages within 1 Gyr, masses within 5% and radii within 10% of the solar values. Their average [X/H] abundances are also consistent with the solar pattern within 2 uncertainties: -0.063 0.037 dex (KOI-444), 0.000 0.057 dex (KOI-3232), 0.036 0.089 dex (KOI-4383) and 0.050 0.069 dex (KOI-4400).
@@
We estimated their chromospheric activity levels through the log(R) index calculated following the methodology described in lt25. The derived values are: -4.937 (KOI-444), -4.646 (KOI-3232), -4.607 (KOI-4383) and -4.851 (KOI-4400). Thus, KOI-444 and KOI-4400 have chromospheric activity levels that are compatible with the solar values (e.g., -4.937 from noyes84, -4.906 from mh08 and -5.021 from lorenzo-oliveira18). They are also consistent with the values determined by isaacson24 (-4.981 for KOI-444 and -5.131 for KOI-4400) and place these twins in the regime of inactive stars, log(R) -4.75, as defined by henry96. All the parameters discussed above make KOI-444 and KOI-4400 the two best solar twins in our sample. The former hosts a warm sub-Neptune, while the latter hosts a hot Earth-sized planet. To the best of our knowledge, this is the first time that these two solar twins are identified in the literature with such a detailed characterization.
@@
To evaluate the slopes for the subsample of solar twins, we performed a weighted least squares fit to the average abundances of all 25 stars discussed above (using the standard deviations as the weights). As before, we use only [X/Fe] abundances to minimize the effects of atomic diffusion (e.g., dotter17; bedell18). We found a positive slope of (13 4) dex K-1, with -value = 3.165 and -value = 0.009. The fit obtained for this sample is shown in the upper right panel of Figure 13 and the Spearman and Pearson tests both confirm that the correlation is not significant (-value 0.001). The null or positive slope suggests that the Sun is similar or refractory depleted relative to our sample of solar twins, which all have planets in this case. In fact, only one of the 25 twins has a negative slope.
@@
If we use the abundances corrected for age trends instead, the correlation is still not significant according to the Spearman and Pearson tests, although the slope is (5 2) dex K-1, which is lower than before but still different from zero within the uncertainty. Another interesting difference is that now 18 out of the 25 solar twins in our sample (72%) have positive slopes and consequently the Sun is more depleted in refractory elements relative to them, a value that is consistent with the range of previous estimates from the literature: 80% from bedell18, 70% from nibauer21, 87% from rampalli24 and 79 - 95% from rampalli25. One main difference between these previous studies and ours is that all solar twins in our sample host exoplanets, making it unlikely that the chemical signature found in previous studies for the Sun may be caused by planet formation.
@@
As discussed above, there is one star among the solar twins that hosts only a single Earth-sized planet (KOI-4400) and its slope is (29 6) dex K-1, with -value = 5.069 and -value = 0.0004. The correlation is significant according only to the Pearson test. Using the [X/Fe] abundances corrected for age trends instead, both tests return -value 0.001. Based on its slope of elemental abundances versus condensation temperature, this star is similar or significantly more enhanced in refractories relative to the Sun, although it hosts an Earth-sized planet.
@@
Since it seems that the slope in the case of solar twins is mainly caused by the three elements with K (Na, Mn and Cu; see Figure 13), we removed them from the computation of the slopes determined from the average abundances and repeated the tests above. Now, the slope is consistent with zero for the solar twins as well as for KOI-4400 and also for the other five subsamples shown in Figure 13. In summary, all our results support the conclusion that the depletion of refractory elements in the Sun is not caused by the presence of planets and/or the system’s architecture.
@@ V Conclusions
We determined chemical abundances for 13 elements and 16 species (Na, Mg, Al, Si, Ca, Sc I, Sc II, Ti I, Ti II, V, Cr I, Cr II, Mn, Co, Ni and Cu) for a sample of 561 stars with planets detected by the Kepler mission (borucki10). We observed that the abundances for all elements are higher than the solar values (with a median of 0.06 dex), possibly reflecting the higher metallicities of the stars. Although the typical S/N of the spectra is 60, we obtained overall precise abundances and the largest median uncertainty is 0.053 for Cr II.
@@
We analyzed the abundances [X/H] as a function of planetary radii as well as orbital period and observed that stars hosting large planets are statistically enriched for most elements relative to those hosting small planets. A similar result was not obtained though when we used [X/Fe] abundances, revealing that the elements are simply following the underlying behavior observed for Fe in ghezzi21.
@@
Going from individual planets to planetary systems, we found statistically significant differences for Al, Si, Sc, Ti, Cr, Fe and Co when comparing the [X/H] abundances for stars with only small and only large planets. However, this result was not confirmed when we adopted [X/Fe] abundances (corrected or not for trends with age) instead, which means that it was probably caused by the well-known planet-metallicity correlation for giant planets (e.g., gonzalez97; fischer05; ghezzi18).
@@
For Al, Si, Ca, Fe, Co and Ni, we also found that the abundances [X/H] for systems with only hot planets are statistically larger than those of systems with only warm planets, a result that was previously found for iron and a few other elements in the literature (e.g., mulders16; petigura18; wilson18; wilson22). Once again, we did not confirm this result with [X/Fe] abundances, suggesting that the other elements simply reflect the higher iron content.
@@
Focusing on systems with only small planets, there are statistically significant differences between Small Hot and Small Warm systems for the abundances [X/H] of Si, Fe and Co. None of the other multiple comparisons we performed yielded statistically significant results. The analysis of groups of elements yielded consistent significant differences between systems with only small and only large planets for the refractories, lithophiles and siderophiles. For the comparisons involving [/Fe] and the elemental ratios Fe/Si and Mg/Si, we found no significant differences.
@@
We analyzed the [Fe/H] distributions for low- and high-[/Fe] stars hosting large planets and found that the minimum metallicity is higher for the former. This result suggests that the -elements are able to partially compensate lower metallicities to allow the formation of giant planets (e.g., adibekyan12a; adibekyan12b). However, the iron abundance still limits this process since the minimum metallicity of the high-[/Fe] systems with small planets is much lower than their counterparts with large planets.
@@
We investigated the slopes of the relation between [X/Fe] abundances and the condensation temperatures of the elements and observed that our stars exhibit a variety of slopes, in agreement with previous studies in the literature (e.g., rampalli24). Multiple tests did not yield statistically significant differences between the distributions of slopes for distinct planetary systems. To the best of our knowledge, this is the largest sample of planet hosting stars for which the study of slopes was performed.
@@
The analysis of 25 solar twins in our sample confirmed that the Sun is refractory depleted relative to 76% of them, a percentage that is consistent with previous works (e.g., bedell18; nibauer21; rampalli24; rampalli25). We also showed that the Sun has at least the same amount of refractory elements or is more depleted than stars hosting super-Earths or Earth-sized planets.
@@
We have identified the stars KOI-444 and KOI-4400 as the best solar twins in our sample since all their atmospheric and evolutionary parameters, abundances and activity levels are consistent with those of the Sun within the typical uncertainties. To the best of our knowledge, this is the first time that both of them are presented with such a thorough characterization. The solar twin KOI-4400 hosts a single Earth-sized planet and is more depleted in refractory elements or has an abundance pattern similar to the Sun.
@@
In summary, our results reveal that all possible chemical signatures found for the planet-hosting stars are caused by the underlying behavior of iron. Moreover, we did not confirm the hypothesis that the depletion of refractory elements in the Sun could be caused by the formation of terrestrial or giant planets. Although we used the largest sample of planet-hosting stars to date in such studies, the spectra have lower S/N values in general. Therefore, we highlight the importance of continuing the ongoing efforts to obtain high-quality spectra and determine more precise physical parameters and chemical abundances for stars with planetary systems in order to fully understand their formation, properties and architectures.

### v2

@@ IV.4.2 Solar twins
Focusing now on host stars with parameters very similar to the solar values, previous works in the literature have found that the Sun is depleted in refractory elements relative to solar twins (e.g., Meléndez et al., 2009; Bedell et al., 2018; Rampalli et al., 2024; Sun et al., 2025b, a). If we assume here that solar twins are stars that have stellar parameters within the ranges 5677 K 5877 K, 4.34 4.54 and -0.10 dex [Fe/H] 0.10 dex, there are 25 solar twins in our sample. Among these systems, we have 6 SE, 11 SN, 6 SE + SN, 1 SN + SS, and 1 JP.
@@
It is worth noting that four of these stars (KOI-444, KOI-3232, KOI-4383, and KOI-4400) have ages within 1 Gyr, masses within 5%, and radii within 10% of the solar values. Their average [X/H] abundances are also consistent with the solar pattern within 2 uncertainties: -0.063 0.037 dex (KOI-444), 0.000 0.057 dex (KOI-3232), 0.036 0.089 dex (KOI-4383), and 0.050 0.069 dex (KOI-4400).
@@
We estimated their chromospheric activity levels through the log(R) index calculated following the methodology described in Loaiza-Tacuri et al. (2025). The derived values are -4.937 (KOI-444), -4.646 (KOI-3232), -4.607 (KOI-4383) and -4.851 (KOI-4400). Thus, KOI-444 and KOI-4400 have chromospheric activity levels that are compatible with the solar values (e.g., -4.937 from Noyes et al. 1984; -4.906 from Mamajek and Hillenbrand 2008; and -5.021 from Lorenzo-Oliveira et al. 2018). They are also consistent with the values determined by Isaacson et al. (2024) (-4.981 for KOI-444 and -5.131 for KOI-4400) and place these twins in the regime of inactive stars, log(R) -4.75, as defined by Henry et al. (1996). All the parameters discussed above make KOI-444 and KOI-4400 the two best solar twins in our sample. The former hosts a warm sub-Neptune, while the latter hosts a hot Earth-sized planet. To the best of our knowledge, this is the first time that these two solar twins are identified in the literature with such a detailed characterization.
@@
To evaluate the slopes for the subsample of solar twins, we performed a WLS fit to the average abundances of all 25 stars discussed above (using the standard deviations as the weights). As before, we use only [X/Fe] abundances to minimize the effects of atomic diffusion (e.g., Dotter et al., 2017; Bedell et al., 2018). We found a positive slope of (13 4) dex K-1, with -value = 3.165 and -value = 0.009. The fit obtained for this sample is shown in the upper right panel of Figure 13, and the Spearman and Pearson tests both confirm that the correlation is not significant (-value 0.001). The null or positive slope suggests that the Sun is similar or refractory depleted relative to our sample of solar twins, which all have planets in this case. In fact, only 1 of the 25 twins has a negative slope.
@@
If we use the abundances corrected for age trends instead, the correlation is still not significant according to the Spearman and Pearson tests, although the slope is (5 2) dex K-1, which is lower than before but still different from zero within the uncertainty. Another interesting difference is that now 18 out of the 25 solar twins in our sample (72%) have positive slopes, and consequently, the Sun is more depleted in refractory elements relative to them, a value that is consistent with the range of previous estimates from the literature: 80% from Bedell et al. (2018), 70% from Nibauer et al. (2021), 87% from Rampalli et al. (2024) and 79 - 95% from Rampalli et al. (2025). One main difference between these previous studies and ours is that all solar twins in our sample host exoplanets, making it unlikely that the chemical signature found in previous studies for the Sun may be caused by planet formation.
@@
As discussed above, there is one star among the solar twins that hosts only a single Earth-sized planet (KOI-4400), and its slope is (29 6) dex K-1, with -value = 5.069 and -value = 0.0004. The correlation is significant according only to the Pearson test. Using the [X/Fe] abundances corrected for age trends instead, both tests return -value 0.001. Based on its slope of elemental abundances versus condensation temperature, this star is similar or significantly more enhanced in refractories relative to the Sun, although it hosts an Earth-sized planet.
@@
Since it seems that the slope in the case of solar twins is mainly caused by the three elements with K (Na, Mn, and Cu; see Figure 13), we removed them from the computation of the slopes determined from the average abundances and repeated the tests above. Now, the slope is consistent with zero for the solar twins as well as for KOI-4400 and also for the other five subsamples shown in Figure 13. In summary, all our results support the conclusion that the depletion of refractory elements in the Sun is not caused by the presence of planets and/or the system’s architecture.
@@ V Conclusions
We determined chemical abundances for 13 elements and 16 species (Na, Mg, Al, Si, Ca, Sc I, Sc II, Ti I, Ti II, V, Cr I, Cr II, Mn, Co, Ni, and Cu) for a sample of 561 stars with planets detected by the Kepler mission (Borucki et al., 2010). We observed that the abundances for all elements are higher than the solar values (with a median of 0.06 dex), possibly reflecting the higher metallicities of the stars. Although the typical S/N of the spectra is 60, we obtained overall precise abundances, and the largest median uncertainty is 0.053 for Cr II.
@@
We analyzed the abundances [X/H] as a function of planetary radii as well as orbital period and observed that stars hosting large planets are statistically enriched for most elements relative to those hosting small planets. A similar result was not obtained, though, when we used [X/Fe] abundances, revealing that the elements are simply following the underlying behavior observed for Fe in Ghezzi et al. (2021).
@@
Going from individual planets to planetary systems, we found statistically significant differences for Al, Si, Sc, Ti, Cr, Fe, and Co when comparing the [X/H] abundances for stars with only small and only large planets. However, this result was not confirmed when we adopted [X/Fe] abundances (corrected or not for trends with age) instead, which means that it was probably caused by the well-known planet-metallicity correlation for giant planets (e.g., Gonzalez, 1997; Fischer and Valenti, 2005; Ghezzi et al., 2018).
@@
For Al, Si, Ca, Fe, Co, and Ni, we also found that the abundances [X/H] for systems with only hot planets are statistically larger than those of systems with only warm planets, a result that was previously found for iron and a few other elements in the literature (e.g., Mulders et al., 2016; Petigura et al., 2018; Wilson et al., 2018, 2022). Once again, we did not confirm this result with [X/Fe] abundances, suggesting that the other elements simply reflect the higher iron content.
@@
Focusing on systems with only small planets, there are statistically significant differences between Small Hot and Small Warm systems for the abundances [X/H] of Si, Fe, and Co. None of the other multiple comparisons we performed yielded statistically significant results. The analysis of groups of elements yielded consistent, significant differences between systems with only small and only large planets for the refractories, lithophiles and siderophiles. For the comparisons involving [/Fe] and the elemental ratios Fe/Si and Mg/Si, we found no significant differences.
@@
We analyzed the [Fe/H] distributions for low- and high-[/Fe] stars hosting large planets and found that the minimum metallicity is higher for the former. This result suggests that the -elements are able to partially compensate for lower metallicities to allow the formation of giant planets (e.g., Adibekyan et al., 2012b, a). However, the iron abundance still limits this process since the minimum metallicity of the high-[/Fe] systems with small planets is much lower than their counterparts with large planets.
@@
We investigated the slopes of the relation between [X/Fe] abundances and the condensation temperatures of the elements and observed that our stars exhibit a variety of slopes, in agreement with previous studies in the literature (e.g., Rampalli et al., 2024). Multiple tests did not yield statistically significant differences between the distributions of slopes for distinct planetary systems. To the best of our knowledge, this is the largest sample of planet-hosting stars for which the study of slopes was performed.
@@
The analysis of 25 solar twins in our sample confirmed that the Sun is refractory depleted relative to 72% of them, a percentage that is consistent with previous works (e.g., Bedell et al., 2018; Nibauer et al., 2021; Rampalli et al., 2024, 2025). We also showed that the Sun has at least the same amount of refractory elements or is more depleted than stars hosting super-Earths or Earth-sized planets.
@@
We have identified the stars KOI-444 and KOI-4400 as the best solar twins in our sample since all their atmospheric and evolutionary parameters, abundances, and activity levels are consistent with those of the Sun within the typical uncertainties. To the best of our knowledge, this is the first time that both of them are presented with such a thorough characterization. The solar twin KOI-4400 hosts a single Earth-sized planet and is more depleted in refractory elements or has an abundance pattern similar to the Sun.
@@
In summary, our results reveal that all possible chemical signatures found for the planet-hosting stars are caused by the underlying behavior of iron. Moreover, we did not confirm the hypothesis that the depletion of refractory elements in the Sun could be caused by the formation of terrestrial or giant planets. Although we used the largest sample of planet-hosting stars to date in such studies, the spectra have lower S/N values in general. Therefore, we highlight the importance of continuing the ongoing efforts to obtain high-quality spectra and determine more precise physical parameters and chemical abundances for stars with planetary systems in order to fully understand their formation, properties, and architectures.